Find the equation of the line AB which passes through (3, 1) and is parallel to 4(y+6)=x

Answer by Alan3354(69443)

Find the equation of the line AB which passes through (3, 1) and is parallel to 4(y+6)=x
----------
Find the slope of the line.
Parallel lines have the same slope.
============
Use y - y1 = m*(x - x1) where (x1,y1) is the point and m = slope